A  cylindrical gas tank 1 m long, inside diameter of 20 cm, is evacuated and then filled with carbon dioxide gas at 25°C. To what pressure should it be charged if there should be 1.2 kg of carbon dioxide?



1
Expert's answer
2021-11-22T10:14:39-0500

"V=l\cdot A=l\cdot\pi d^2/4=1\cdot3.14\cdot0.2^2/4=0.1256\ (m^3)"


"pV=mRT/M\to p=mRT/(MV)="


"p=1.2\cdot8.31\cdot298/(0.048\cdot0.1256)\approx493\ (kPa)" . Answer